I is between H and J. HJ = 20x-32, HI 7x+32, and IJ =12x+32. Find HJ
1
Expert's answer
2012-10-01T09:56:27-0400
HJ=HI+IJ=7x+32+12x+32=20x-32
19x+64=20x-32
x=96=3*32
HJ=20*3*32-32=59*32=1888